Page tree
Skip to end of metadata
Go to start of metadata

Attendees

Apologies

Agenda

      • SATA disk, or other problem with SoloHI 00.02 series of VMs.
      • Keep old log files for how long... (PHI currently planning on 6 months)
      • VM status - Engineering.
      • VM product status.
      • AOB

Discussion items

ItemNotes

VM disk issues (SATA?)

SoloHI confirmed that their series 02 VMs were indeed using a SATA disk, and that they would change this. Andrew Lanagan mentioned that their system was relatively minimal based on Ubuntu Server.

Log retention

RC raised the issue of how long instrument VMs were retaining their rotated log files. He encouraged teams to consider how long they thought useful.

VM Status - engineering.

MP said that he was currently implementing the handling of lightcurves. He hoped to have that ready by the end of February. DW asked if he was working with flight data. Maurizio said that yes - with data from November. He believes with the new structure, after the lightcurve header change. DW to give feedback to MP on the current headers.

VE mentioned that MAG were not able to update the VM yet.

XB raised the issue of the slow response of the RPW VM to the first processing request. RC mentioned that this phenomenon is also seen with other VMs - the first request takes longer to process - although the RPW case was just a bit more pronounced, leading to the tests timing out. This had happened just occasionally with one of two other instruments' VM but RC didn't remember which. Since it is not the plan to restart the VMs regularly, the issue is not major.

Shane asked that the SOC check again whether STIX log rotation is taking place. Since the meeting RC has looked and at the time of writing (2021-03-25) there has been no rotation and the log is 653.2 MB.

VM product status

AW mentioned that the RPW filename issue had been fixed. In addition the EDP duplicates had been fixed.
Re SPICE DW mentioned that he had spotted one small issue with the header - a RAST_NUM keyword.

Terje mentioned that he hasn't changed the LLDPDD yet, but has made a few changes in the VM, including one only interesting for experts.

Dave said the the SPICE LLDPDD looks good.

Dave then mentioned that with STIX there were a few smallish issues, the most notable was filenames with underscores (where there should be dashes??). He was planning to send an email on this topic on the day of the meeting. Shane said he should be able to quite quickly create and upload a new VM.

DW mentioned that he had looked at the data coming from Metis VM 00.06.05 corresponding to Cape telemetry. There were quite a few issues with the metadata. He asked if this has been changed. MP said he was working on that right then.

AoB

Shane asked if it was possible to get the LLPH output files from the SOC. RC said he was happy to provide them occasionally. And DW pointed put that they would be available from the archive after sending them there was enabled.

It was agreed the next telecon should be on the 24th of March.

Preparatory Actions

New Actions

None.